Calculate the pH of a 0.025 M solution of CH.NHCI (Show your work 2 pts). C5H5NHCl


1
Expert's answer
2021-03-15T09:06:17-0400

"K_a=\frac{K_w} {K_b} =\frac{1×10^{-14}}{1.7×10^{-9}}"


"=5.88×10^{-6}=\frac{x^{2}} {0.025-x}=\frac{x^{2}}{0.025}"


"x=1.21\u00d710^{-2}M=[H^{+}]"


"pH=-log(1.21\u00d710^{-2})=1.92"